# On face angles of tetrahedra with a given base

### v2

@@ 5. Properties of the map $F$ on the plane $ABC$
Proof. Since $D\in\Pi^{0}\setminus\{A,B,C\}$, then the volume of the tetrahedron $ABCD$ is zero. Hence, $01+2\cos\overline{\alpha}\cdot\cos\overline{\beta}\cdot\cos\overline{\gamma}-\cos^{2}\overline{\alpha}-\cos^{2}\overline{\beta}-\cos^{2}\overline{\gamma}=$ 0 by (5).
@@
$\Pi^{0}=\mathbb{B}_{-}\bigcup S\bigcup\mathbb{B}_{+},$ (24)
@@
$\widetilde{A}$ ::=:= $\Bigl{(}-\cos\angle BAC,\cos\angle ABC,\cos\angle ACB\Bigl{)},$
$\widetilde{B}$ ::=:= $\Bigl{(}\cos\angle BAC,-\cos\angle ABC,\cos\angle ACB\Bigl{)},$ (25)
$\widetilde{C}$ ::=:= $\Bigl{(}\cos\angle BAC,\cos\angle ABC,-\cos\angle ACB\Bigl{)}.$
@@
The map $F$ is locally constant and takes only three values on the set $S\setminus\{A,B,C\}$. Moreover, if $D$ is from the arc $BC$, $AC$, or $AB$, then $F(D)$ is equal to $\widetilde{A}$, $\widetilde{B}$, or $\widetilde{C}$, respectively.
@@
Let us suppose that there are $D_{1},D_{2}\in\mathbb{B}_{-}$ such that $D_{1}\neq D_{2}$ and $F(D_{1})=F(D_{2})$. The segment $[D_{1},D_{2}]$ cannot intersect all sides of the triangle $ABC$. Suppose that it does not intersect the side $BC$. Then $[D_{1},D_{2}]$ should intersect two other sides of the triangle $ABC$. Indeed, if it does not intersect, say the side $AB$, then both points $D_{1},D_{2}$ are on a circle $\mathcal{C}_{1}$ with the chord $AB$ as well as are on a circle $\mathcal{C}_{2}$ with the chord $BC$. Hence, $D_{1},D_{2},B\in\mathcal{C}_{1}\bigcap\mathcal{C}_{2}$ that impossible (an intersection of two circles has at most two points). Now, we may suppose without loss of generality that $D_{1}$ and $C$ are in different half-planes with respect the straight line $AB$ and $[D_{1},D_{2}]$ intersect the sides $AB$ and $AC$. Hence, we have a convex pentagon $AD_{1}BCD_{2}$, which implies $\angle AD_{1}C<\angle AD_{1}B$ and $\angle AD_{2}C>\angle AD_{2}B$. But the condition $F(D_{1})=F(D_{2})$ implies $\angle AD_{1}C=\angle AD_{2}C$ and $\angle AD_{1}B=\angle AD_{2}B$, which is incompatible with the previous inequalities. Hence, $F$ is injective on the set $\mathbb{B}_{-}$.
@@
$-\cos\angle CBA<-\cos\angle ADC=\cos\angle BDC<\cos\angle CAB.$
Therefore, the image of the interval $(A,B)$ under the map $F$ is the interval between the points $(\cos\angle CAB,-\cos\angle CAB,-1)$ and $(-\cos\angle ABC,\cos\angle ABC,-1)$, which is a subset of the interval $I_{1}$ with the endpoints $(-1,1,-1)$ $(1,-1,-1)$.
The same idea implies that the image of the interval $(B,C)$ under the map $F$ is the interval between the endpoints $(-1,\cos\angle ACB,-\cos\angle ACB)$ and $(-1,-\cos\angle ABC,\cos\angle ABC)$, which is a subset of the interval $I_{2}$ with the endpoints $(-1,-1,1)$ $(-1,1,-1)$.
Finally, similar arguments show that the image of the interval $(A,C)$ under the map $F$ is the interval between the point $(-\cos\angle ABC,-1,\cos\angle ABC)$ and the point $(\cos\angle CAB,-1,-\cos\angle CAB)$, which is a subset of the interval $I_{3}$ with the endpoints $(1,-1,-1)$ $(-1,-1,1)$.
@@
Figure 7. Intersection of the toroid $\mathbb{T}_{CB}^{\alpha}$ with the plane $ABC$
@@
It is easy to check the following observation: if $\alpha<\min\{\angle ACB,\angle ABC\}$, then the curve $\gamma_{1}$ has its endpoints on the ray $AC$ and on the ray $AB$; $\angle CDA$ can take any value in $[0,\alpha]$ for points $D\in\gamma_{1}$; and $\angle BDA=\alpha-\angle CDA$, see Fig. 7. Therefore, any triple $(\alpha,\alpha_{1},\alpha_{2})$ such that $\alpha_{1}\geq 0$, $\alpha_{2}\geq 0$, and $\alpha=\alpha_{1}+\alpha_{2}$, can be realized as $\left(\angle BDC,\angle CDA,\angle BDA\right)$ for some $D\in\gamma_{1}$.
@@
${\mathbb{G}_{1}^{0}\,(\mathbb{G}_{2}^{0})}$ ::=:= $\{D\in\mathbb{G}_{1}\,(\mathbb{G}_{2})\,|\,\angle BDC\in\left(0,\min\{\angle ACB,\angle ABC\}\right)\},$
${\mathbb{G}_{3}^{0}\,(\mathbb{G}_{4}^{0})}$ ::=:= $\{D\in\mathbb{G}_{3}\,(\mathbb{G}_{4})\,|\,\angle ADC\in\left(0,\min\{\angle BAC,\angle BCA\}\right)\},$
${\mathbb{G}_{5}^{0}\,(\mathbb{G}_{6}^{0})}$ ::=:= $\{D\in\mathbb{G}_{5}\,(\mathbb{G}_{6})\,|\,\angle ADB\in\left(0,\min\{\angle CAB,\angle CBA\}\right)\}.$
@@
$F({\mathbb{G}_{2}^{0}})\subset F({\mathbb{G}_{1}^{0}}),\quad F({\mathbb{G}_{4}^{0}})\subset F({\mathbb{G}_{3}^{0}}),\quad F({\mathbb{G}_{6}^{0}})\subset F({\mathbb{G}_{5}^{0}}).$ (26)
@@
$[\infty,D_{1})\leftrightarrow[\infty,D_{2}),\quad[\infty,D_{3})\leftrightarrow[\infty,D_{4}),\quad[\infty,D_{5})\leftrightarrow[\infty,D_{6}).$
@@
The map $F$ is injective in the interior of any of the angular domains $ABC$, $BCA$, $CAB$.
Proof. It suffices to consider only the angular domain $ABC$ ($\mathbb{G}_{0}\bigcup\mathbb{G}_{3}$). Let us suppose that there are $D_{1}\neq D_{2}$ in this angular domain such that $F(D_{1})=F(D_{2})$, Then $D_{1}$ and $D_{2}$ lie on a circle $C_{1}$ with chord $[AB]$ ($\angle BD_{1}A=\angle BD_{2}A$). By similar reasons $D_{1}$ and $D_{2}$ lie on a circle $C_{2}$ with chord $[BC]$. Therefore, $D_{1},D_{2}\in C_{1}\bigcap C_{2}$. On the other hand, $B\in C_{1}\bigcap C_{2}$. Since the intersection of two distinct circles has at most two points, we get the proposition.
@@
Let us consider $\mathfrak{m}:=\max\{\cos\angle BAC,\cos\angle ABC,\cos\angle BCA\}$ for a given base $ABC$. If $(1,1,1)\neq(\tilde{x},\tilde{y},\tilde{z})\in\mathbb{BP}$ and the inequality $\min\{\tilde{x},\tilde{y},\tilde{z}\}>\mathfrak{m}$ holds, then there exists $D\in\Pi^{0}\setminus\{A,B,C\}$ such that $F(D)=(\tilde{x},\tilde{y},\tilde{z})$.
Figure 9. Intersection of the toroid $\mathbb{T}_{AB}^{\,\overline{\gamma}}$ with a special angular domain on the plane $ABC$
Proof. It is clear that $\mathfrak{m}\geq 1/2=\cos(\pi/3)$. Without loss of generality we assume that $\tilde{x}\geq\tilde{y}\geq\tilde{z}>\mathfrak{m}\geq 1/2$. Let us consider the numbers $\overline{\alpha}:=\arccos(\tilde{x})$, $\overline{\beta}:=\arccos(\tilde{y})$, and $\overline{\gamma}:=\arccos(\tilde{z})$. Obviously, $30\leq\overline{\alpha}\leq\overline{\beta}\leq\overline{\gamma}<\pi/$ 3. Using Lemma 2, we can suppose that
@@
$F(D)=(\cos\angle BDC,\cos\angle ADC,\cos\angle BDA)=(\tilde{x},\tilde{y},\tilde{z}).$
@@ 6. On the limit points of the image $F(\operatorname{GT})$
For a given nontrivial vector $\eta\in\mathbb{R}^{3}$, we consider the points $D=B+t\cdot\eta$, $t\in\mathbb{R}$. If the third coordinate of $\eta$ is not 00, then we consider $t$ only with the same sign as this coordinate has. It is easy to check that
@@
$(\cos\overline{\alpha},\cos\overline{\beta},\cos\overline{\gamma})\rightarrow\left(-\frac{(\tau_{2},\eta_{h})}{\|\eta\|},\cos\angle ABC,\frac{(\tau_{1},\eta_{h})}{\|\eta\|}\right).$ (27)
@@
$\operatorname{Lim}(B)=\left\{\bigl{(}x,y_{0},z\bigr{)}\in\mathbb{R}^{3}\,\left|\,x^{2}+z^{2}-2y_{0}xz\leq 1-y_{0}^{2},\,y_{0}=\cos\angle ABC\right.\right\}.$
@@
$\operatorname{Lim}(A)=\left\{\bigl{(}x_{0},y,z\bigr{)}\in\mathbb{R}^{3}\,\left|\,y^{2}+z^{2}-2x_{0}yz\leq 1-x_{0}^{2},\,x_{0}=\cos\angle BAC\right.\right\},$
$\operatorname{Lim}(B)=\left\{\bigl{(}x,y_{0},z\bigr{)}\in\mathbb{R}^{3}\,\left|\,x^{2}+z^{2}-2y_{0}xz\leq 1-y_{0}^{2},\,y_{0}=\cos\angle ABC\right.\right\},$
$\operatorname{Lim}(C)=\left\{\bigl{(}x,y,z_{0}\bigr{)}\in\mathbb{R}^{3}\,\left|\,x^{2}+y^{2}-2z_{0}xy\leq 1-z_{0}^{2},\,z_{0}=\cos\angle ACB\right.\right\}.$
@@
$01-2\cos\angle BAC\cdot\cos\angle ABC\cdot\cos\angle ACB-\cos^{2}\angle BAC-\cos^{2}\angle ABC-\cos^{2}\angle ACB=$ 0 .
@@
[…]
@@
$E_{A}=\left\{\bigl{(}x_{0},y,z\bigr{)}\in\mathbb{R}^{3}\,\left|\,y^{2}+z^{2}-2x_{0}yz=1-x_{0}^{2},\,x_{0}=\cos\angle BAC\right.\right\},$
$E_{B}=\left\{\bigl{(}x,y_{0},z\bigr{)}\in\mathbb{R}^{3}\,\left|\,x^{2}+z^{2}-2y_{0}xz=1-y_{0}^{2},\,y_{0}=\cos\angle ABC\right.\right\},$
$E_{C}=\left\{\bigl{(}x,y,z_{0}\bigr{)}\in\mathbb{R}^{3}\,\left|\,x^{2}+y^{2}-2z_{0}xy=1-z_{0}^{2},\,z_{0}=\cos\angle ACB\right.\right\}.$
@@ 7. On the images of the points of $\operatorname{Cyl}$
We know that $F(D)=\widetilde{A}=\Bigl{(}-\cos\angle BAC,\cos\angle ABC,\cos\angle ACB\Bigr{)}$, if the point $D$ lies on the arc $BC$; $F(D)=\widetilde{B}=\Bigl{(}\cos\angle BAC,-\cos\angle ABC,\cos\angle ACB\Bigr{)}$, if the point $D$ lies on the arc $AC$ and $F(D)=\widetilde{C}=\Bigl{(}\cos\angle BAC,\cos\angle ABC,-\cos\angle ACB\Bigr{)}$, if the point $D$ lies on the arc $AB$ (see (5) and Lemma 5). Therefore, as we already know, the function $F$ on the set $S\backslash\{A,B,C\}$ can take only three possible values.
@@
where $\eta_{h}=\lambda\cdot e_{B}$, $\eta_{v}=e_{3}$ (note that $(\tau_{1},\tau_{2})=-\cos\angle ABC$). Therefore, the limit points of the points $F(D)$, where $D\in\operatorname{Cyl}$ and $D\to B$, is a segment with the endpoints $\bigl{(}-(\tau_{2},e_{B}),\cos\angle ABC,(\tau_{1},e_{B})\bigr{)}$ and $\bigl{(}(\tau_{2},e_{B}),\cos\angle ABC,-(\tau_{1},e_{B})\bigr{)}$. On the other hand, if $D$ tends to $B$ along the arc $AB$ ($BC$), then $F(D)=\widetilde{C}$ (respectively, $F(D)=\widetilde{A}$). Hence, $(\tau_{1},e_{B})=\pm\cos\angle ACB$ and $(\tau_{2},e_{B})=\pm\cos\angle BAC$.
@@
$[\widetilde{B},\widetilde{C}]\subset\operatorname{Lim}(A),\quad[\widetilde{A},\widetilde{C}]\subset\operatorname{Lim}(B),\quad[\widetilde{A},\widetilde{B}]\subset\operatorname{Lim}(C),$
@@
If we have a closed curve $\Gamma$, which goes through points in all arcs $BC$, $AC$, and $AB$, then the image $F(\Gamma)$ is closed and goes through the points $\widetilde{A}$, $\widetilde{B}$, and $\widetilde{C}$.
@@
$r^{2}\Bigl{(}F(p,q,r)-(1,1,1)\Bigr{)}\to\bigl{(}\cos(\mu-\nu)-1,\cos(\nu)-1,\cos(\mu)-1\bigr{)}$ (28)
@@
$\frac{(v^{2}+1)(t^{2}+1)r^{2}+(2tv-1)^{2}-1}{\sqrt{v^{2}+1}\sqrt{P_{1}}\sqrt{P_{3}}},\quad\frac{(u^{2}+1)(t^{2}+1)r^{2}+(2tu-1)^{2}-1}{\sqrt{u^{2}+1}\sqrt{P_{1}}\sqrt{P_{2}}}.$
@@
$\frac{16r^{2}(t+v)(t^{2}+2tv-1)}{\sqrt{v^{2}+1}\cdot P_{1}^{3/2}\cdot P_{3}^{3/2}},\quad\frac{16r^{2}(t+u)(t^{2}+2tu-1)}{\sqrt{u^{2}+1}\cdot P_{1}^{3/2}\cdot P_{2}^{3/2}}.$
@@
This equation has the roots $\theta=\frac{1}{3}\bigl{(}\pi(2k+1)-\mu-\nu\bigr{)}$, where $k$ is integer and $k\in\left.\left[\frac{\mu+\nu}{2\pi}-1/2,\frac{\mu+\nu}{2\pi}+5/2\right.\right)$.
@@
It should be noted that $r>0$ and $(t^{2}+3)^{2}-12t^{2}=(t^{2}-3)^{2}\geq 0$. In this case $M_{\,\operatorname{reg}}=\bigl{\{}(\cos\theta,\sin\theta,r)\,|\,\theta\in[0,2\pi)\setminus\{\pi/3,\pi,5\pi/3\},\,r\in(0,\infty)\bigr{\}}$. We note that the images of the points $(1/2,\sqrt{3}/2,r)$, $(-1,0,r)$, $(-1/2,\sqrt{3}/2,r)$ under the map $F$ are non-regular points of the surface $\operatorname{FC}=F(\operatorname{Cyl}\bigcap\Pi^{+})$ for any positive $r$.
@@
$p=\cos\varphi+\eta_{1}\cdot t,\quad q=\sin\varphi+\eta_{2}\cdot t,\quad r=z+\eta_{3}\cdot t$ (30)
@@
$F_{i}\bigl{(}\cos\varphi+\eta_{1}\cdot t,\sin\varphi+\eta_{2}\cdot t,z+\eta_{3}\cdot t\Bigr{)}=F_{i}\bigl{(}\cos\varphi,\sin\varphi,z\bigr{)}+\xi_{i}\cdot t^{2}+o(t^{2}),$ (31)
@@
Moreover, $\xi_{1}^{2}$, $\xi_{2}^{2}$, and $\xi_{3}^{2}$ are rational functions with respect to $z$, such that the degree of the numerator coincides with the degree of the denominator and is equal to $20$, $16$, $16$, respectively.
@@
$\cos\angle BAC\cdot\cos\angle ABC\cdot\cos\angle ACB=-\frac{1+uv}{(1+u^{2})(1+v^{2})}$ (32)
@@
It should be noted that the point $D=(p_{0},q_{0},r_{0})$ in the proof of Proposition 19 is above the de Longchamps point for $\triangle ABC$, which is the reflection of the orthocenter of $\triangle ABC$ about the circumcenter. The see this, note that as vectors, $A+D$, $B+D$, $C+D$ are orthogonal to $B-C$, $C-A$, $A-B$, respectively.

### v3

@@ 5. Properties of the map $F$ on the plane ABCABC
Proof. Since $D\in\Pi^{0}\setminus\{A,B,C\}$, then the volume of the tetrahedron $ABCD$ is zero. Hence, $1+2\cos\overline{\alpha}\cdot\cos\overline{\beta}\cdot\cos\overline{\gamma}-\cos^{2}\overline{\alpha}-\cos^{2}\overline{\beta}-\cos^{2}\overline{\gamma}=0$ by (5).
@@
(24) $\Pi^{0}=\mathbb{B}_{-}\bigcup S\bigcup\mathbb{B}_{+},$
@@
$\widetilde{A}$ $:=$ $\Bigl(-\cos\angle BAC,\cos\angle ABC,\cos\angle ACB\Bigl),$
(25) $\widetilde{B}$ $:=$ $\Bigl(\cos\angle BAC,-\cos\angle ABC,\cos\angle ACB\Bigl),$
$\widetilde{C}$ $:=$ $\Bigl(\cos\angle BAC,\cos\angle ABC,-\cos\angle ACB\Bigl).$
@@
The map $F$ is locally constant and takes only three values on the set $S\setminus\{A,B,C\}$. Moreover, if $D$ is from the arc $BC$, ACAC, or $AB$, then $F(D)$ is equal to $\widetilde{A}$, $\widetilde{B}$, or $\widetilde{C}$, respectively.
@@
Let us suppose that there are $D_{1},D_{2}\in\mathbb{B}_{-}$ such that $D_{1}\neq D_{2}$ and $F(D_{1})=F(D_{2})$. The segment $D_{1},D_{2}]$ cannot intersect all sides of the triangle $ABC$. Suppose that it does not intersect the side $BC$. Then $D_{1},D_{2}]$ should intersect two other sides of the triangle $ABC$. Indeed, if it does not intersect, say the side $AB$, then both points $D_{1},D_{2}$ are on a circle $\mathcal{C}_{1}$ with the chord $AB$ as well as are on a circle $\mathcal{C}_{2}$ with the chord $BC$. Hence, $D_{1},D_{2},B\in\mathcal{C}_{1}\bigcap\mathcal{C}_{2}$ that impossible (an intersection of two circles has at most two points). Now, we may suppose without loss of generality that $D_{1}$ and $C$ are in different half-planes with respect the straight line $AB$ and $D_{1},D_{2}]$ intersect the sides $AB$ and $AC$. Hence, we have a convex pentagon $AD_{1}BCD_{2}$, which implies $\angle AD_{1}C<\angle AD_{1}B$ and $\angle AD_{2}C>\angle AD_{2}B$. But the condition $F(D_{1})=F(D_{2})$ implies $\angle AD_{1}C=\angle AD_{2}C$ and $\angle AD_{1}B=\angle AD_{2}B$, which is incompatible with the previous inequalities. Hence, $F$ is injective on the set $\mathbb{B}_{-}$.
@@
$\cos\angle CBA<-\cos\angle ADC=\cos\angle BDC<\cos\angle CAB.$
Therefore, the image of the interval $(A,B)$ under the map $F$ is the interval between the points $(\cos\angle CAB,-\cos\angle CAB,-1)$ and $(-\cos\angle ABC,\cos\angle ABC,-1)$, which is a subset of the interval $I_{1}$ with the endpoints (1,1,1)(-1,1,-1) и (1,1,1)(1,-1,-1).
The same idea implies that the image of the interval $(B,C)$ under the map $F$ is the interval between the endpoints $(-1,\cos\angle ACB,-\cos\angle ACB)$ and $(-1,-\cos\angle ABC,\cos\angle ABC)$, which is a subset of the interval $I_{2}$ with the endpoints (1,1,1)(-1,-1,1) и (1,1,1)(-1,1,-1).
Finally, similar arguments show that the image of the interval $(A,C)$ under the map $F$ is the interval between the point $(-\cos\angle ABC,-1,\cos\angle ABC)$ and the point $(\cos\angle CAB,-1,-\cos\angle CAB)$, which is a subset of the interval $I_{3}$ with the endpoints (1,1,1)(1,-1,-1) и (1,1,1)(-1,-1,1).
@@
Figure 7. Intersection of the toroid $\mathbb{T}_{CB}^{\alpha}$ with the plane ABCABC
@@
It is easy to check the following observation: if $\alpha<\min\{\angle ACB,\angle ABC\}$, then the curve $\gamma_{1}$ has its endpoints on the ray $AC$ and on the ray $AB$; $\angle CDA$ can take any value in $0,\alpha]$ for points $D\in\gamma_{1}$; and $\angle BDA=\alpha-\angle CDA$, see Fig. 7. Therefore, any triple $(\alpha,\alpha_{1},\alpha_{2})$ such that $\alpha_{1}\geq 0$, $\alpha_{2}\geq 0$, and $\alpha=\alpha_{1}+\alpha_{2}$, can be realized as $\left(\angle BDC,\angle CDA,\angle BDA\right)$ for some $D\in\gamma_{1}$.
@@
${\mathbb{G}_{1}^{0}\,(\mathbb{G}_{2}^{0})}$ $:=$ $\{D\in\mathbb{G}_{1}\,(\mathbb{G}_{2})\,|\,\angle BDC\in\left(0,\min\{\angle ACB,\angle ABC\}\right)\},$
${\mathbb{G}_{3}^{0}\,(\mathbb{G}_{4}^{0})}$ $:=$ $\{D\in\mathbb{G}_{3}\,(\mathbb{G}_{4})\,|\,\angle ADC\in\left(0,\min\{\angle BAC,\angle BCA\}\right)\},$
${\mathbb{G}_{5}^{0}\,(\mathbb{G}_{6}^{0})}$ $:=$ $\{D\in\mathbb{G}_{5}\,(\mathbb{G}_{6})\,|\,\angle ADB\in\left(0,\min\{\angle CAB,\angle CBA\}\right)\}.$
@@
(26) $F({\mathbb{G}_{2}^{0}})\subset F({\mathbb{G}_{1}^{0}}),\quad F({\mathbb{G}_{4}^{0}})\subset F({\mathbb{G}_{3}^{0}}),\quad F({\mathbb{G}_{6}^{0}})\subset F({\mathbb{G}_{5}^{0}}).$
@@
$\infty,D_{1})\leftrightarrow[\infty,D_{2}),\quad[\infty,D_{3})\leftrightarrow[\infty,D_{4}),\quad[\infty,D_{5})\leftrightarrow[\infty,D_{6}).$
@@
The map $F$ is injective in the interior of any of the angular domains $ABC$, BCABCA, CABCAB.
Proof. It suffices to consider only the angular domain ABCABC ($\mathbb{G}_{0}\bigcup\mathbb{G}_{3}$). Let us suppose that there are $D_{1}\neq D_{2}$ in this angular domain such that $F(D_{1})=F(D_{2})$, Then $D_{1}$ and $D_{2}$ lie on a circle $C_{1}$ with chord $[AB]$ ($\angle BD_{1}A=\angle BD_{2}A$). By similar reasons $D_{1}$ and $D_{2}$ lie on a circle $C_{2}$ with chord $[BC]$. Therefore, $D_{1},D_{2}\in C_{1}\bigcap C_{2}$. On the other hand, $B\in C_{1}\bigcap C_{2}$. Since the intersection of two distinct circles has at most two points, we get the proposition.
@@
[…]
Figure 9. Intersection of the toroid $\mathbb{T}_{AB}^{\,\overline{\gamma}}$ with a special angular domain on the plane ABCABC
[…]
@@
[…]
@@ 6. On the limit points of the image $F(\operatorname{GT})$
For a given nontrivial vector $\eta\in\mathbb{R}^{3}$, we consider the points $D=B+t\cdot\eta$, $t\in\mathbb{R}$. If the third coordinate of $\eta$ is not 0, then we consider $t$ only with the same sign as this coordinate has. It is easy to check that
@@
(27) $(\cos\overline{\alpha},\cos\overline{\beta},\cos\overline{\gamma})\rightarrow\left(-\frac{(\tau_{2},\eta_{h})}{\|\eta\|},\cos\angle ABC,\frac{(\tau_{1},\eta_{h})}{\|\eta\|}\right).$
@@
[…]
@@
[…]
@@
$1-2\cos\angle BAC\cdot\cos\angle ABC\cdot\cos\angle ACB-\cos^{2}\angle BAC-\cos^{2}\angle ABC-\cos^{2}\angle ACB=0.$
@@
The closure of $F\Bigl(\Pi^{0}\setminus\{A,B,C\}\Bigr)$ is a subset of the ‘‘pillowcase’’ $\mathbb{BP}=\left\{(x,y,z)\in[-1,1]^{3}\,|\,1+2xyz-x^{2}-y^{2}-z^{2}=0\right\}$ and has the following form:
@@
[…]
@@ 7. On the images of the points of $\operatorname{Cyl}$
[…]
@@
where $\eta_{h}=\lambda\cdot e_{B}$, $\eta_{v}=e_{3}$ (note that $(\tau_{1},\tau_{2})=-\cos\angle ABC$). Therefore, the limit points of the points $F(D)$, where $D\in\operatorname{Cyl}$ and $D\to B$, is a segment with the endpoints $\bigl(-(\tau_{2},e_{B}),\cos\angle ABC,(\tau_{1},e_{B})\bigr)$ and $\bigl((\tau_{2},e_{B}),\cos\angle ABC,-(\tau_{1},e_{B})\bigr)$. On the other hand, if $D$ tends to $B$ along the arc ABAB (BCBC), then $F(D)=\widetilde{C}$ (respectively, $F(D)=\widetilde{A}$). Hence, $(\tau_{1},e_{B})=\pm\cos\angle ACB$ and $(\tau_{2},e_{B})=\pm\cos\angle BAC$.
@@
$\widetilde{B},\widetilde{C}]\subset\operatorname{Lim}(A),\quad[\widetilde{A},\widetilde{C}]\subset\operatorname{Lim}(B),\quad[\widetilde{A},\widetilde{B}]\subset\operatorname{Lim}(C),$
@@
If we have a closed curve $\Gamma$, which goes through points in all arcs $BC$, ACAC, and $AB$, then the image $F(\Gamma)$ is closed and goes through the points $\widetilde{A}$, $\widetilde{B}$, and $\widetilde{C}$.
@@
(28) $r^{2}\Bigl(F(p,q,r)-(1,1,1)\Bigr)\to\bigl(\cos(\mu-\nu)-1,\cos(\nu)-1,\cos(\mu)-1\bigr)$
@@
$\frac{(v^{2}+1)(t^{2}+1)r^{2}+(2tv-1)^{2}-1}{\sqrt{v^{2}+1}\sqrt{P_{1}}\sqrt{P_{3}}},\qquad\frac{(u^{2}+1)(t^{2}+1)r^{2}+(2tu-1)^{2}-1}{\sqrt{u^{2}+1}\sqrt{P_{1}}\sqrt{P_{2}}}.$
@@
$\frac{16r^{2}(t+v)(t^{2}+2tv-1)}{\sqrt{v^{2}+1}\cdot P_{1}^{3/2}\cdot P_{3}^{3/2}},\qquad\frac{16r^{2}(t+u)(t^{2}+2tu-1)}{\sqrt{u^{2}+1}\cdot P_{1}^{3/2}\cdot P_{2}^{3/2}}.$
@@
[…]
@@
It should be noted that $r>0$ and $(t^{2}+3)^{2}-12t^{2}=(t^{2}-3)^{2}\geq 0$. In this case $M_{\,\operatorname{reg}}=\bigl\{(\cos\theta,\sin\theta,r)\,|\,\theta\in[0,2\pi)\setminus\{\pi/3,\pi,5\pi/3\},\,r\in(0,\infty)\bigr\}$. We note that the images of the points $(1/2,\sqrt{3}/2,r)$, (1,0,r)(-1,0,r), $(-1/2,\sqrt{3}/2,r)$ under the map $F$ are non-regular points of the surface $\operatorname{FC}=F(\operatorname{Cyl}\bigcap\Pi^{+})$ for any positive $r$.
@@
(30) $p=\cos\varphi+\eta_{1}\cdot t,\quad q=\sin\varphi+\eta_{2}\cdot t,\quad r=z+\eta_{3}\cdot t$
@@
(31) $F_{i}\bigl(\cos\varphi+\eta_{1}\cdot t,\sin\varphi+\eta_{2}\cdot t,z+\eta_{3}\cdot t\Bigr)=F_{i}\bigl(\cos\varphi,\sin\varphi,z\bigr)+\xi_{i}\cdot t^{2}+o(t^{2}),$
@@
Moreover, $\xi_{1}^{2}$, $\xi_{2}^{2}$, and $\xi_{3}^{2}$ are rational functions with respect to $z$, such that the degree of the numerator coincides with the degree of the denominator and is equal to $20$, 1616, 1616, respectively.
@@
(32) $\cos\angle BAC\cdot\cos\angle ABC\cdot\cos\angle ACB=-\frac{1+uv}{(1+u^{2})(1+v^{2})}$
@@
It should be noted that the point $D=(p_{0},q_{0},r_{0})$ in the proof of Proposition 19 is above the de Longchamps point for $\triangle ABC$, which is the reflection of the orthocenter of $\triangle ABC$ about the circumcenter. The see this, note that as vectors, $A+D$, $B+D$, $C+D$ are orthogonal to BCB-C, CAC-A, ABA-B, respectively.